These reaction times (in tenths of a second) were recorded for group of subjects after each had been given a drug pain.



Drug A
4
7
6
3
4
3

Drug B
9
11
12
8
10
11

Drug C
8
6
7
6
5
7

Formulate the H0, and H1:
C.V. at alpha .01 level of significance
Compute the test value: ANOVA – F ratio
Decision
Interpretation
1
Expert's answer
2020-06-15T19:33:15-0400
"\def\arraystretch{1.5}
 \begin{array}{c:c}
 & Drug\ A & Drug\ B & Drug\ C \\ \hline
 & 4 & 9 & 8 \\
 & 7 & 11 & 6 \\
 & 6 & 12 & 7 \\
 & 3 & 8 & 6 \\
 & 4 & 10 & 5 \\
 & 3 & 11 & 7 \\
 & & & \\
 Sum= & 27 & 61 & 39 \\
 Average= & 4.5 & 10.167 & 6.5 \\
 \sum_ix_{ij}^2= & 135 & 631 & 259 \\
 St.Dev= & 1.643 & 1.472 & 1.049 \\
 SS= & 13.5 & 10.833 & 5.5 \\
 n= & 6 & 6 & 6
\end{array}"

"\bar{x}={1\over n}\sum_ix_i""s^2={1\over n-1}\sum_i(x_i-\bar{x})^2""\bar{x}_A=4.5, s_A^2=2.7,n_A=6"

"\bar{x}_B={61\over 6}, s_B^2={13\over 6},n_B=6"

"\bar{x}_C=6.5, s_C^2=1.1,n_C=6"


The total sample size is "N=18." Therefore, the total degrees of freedom are:


"df_{total}=18-1=17"

Also, the between-groups degrees of freedom are "df_{between}=3-1=2" , and the within-groups degrees of freedom are:


"df_{within}=df_{total}-df_{between}=17-2=15"

Compute the total sum of values and the grand mean. The following is obtained


"\displaystyle\sum_{i,j}x_{ij}=27+61+39=127"

The sum of squared values is


"\displaystyle\sum_{i,j}x_{ij}^2 =135+631+259=1025"

The total sum of squares is computed as follows


"SS_{total}=\displaystyle\sum_{i,j}x_{ij}^2-{1\over N}(\displaystyle\sum_{i,j}x_{ij})^2=1025-{127^2\over 18}=128.944"

The within sum of squares is computed as shown in the calculation below:


"SS_{within}=\displaystyle\sum SS_{within\ groups}=""=13.5+10.833+5.5=29.833"

The between sum of squares is computed directly as shown in the calculation below:

The between sum of squares is computed directly as shown in the calculation below:

Now that sum of squares are computed, we can proceed with computing the mean sum of squares:


"MS_{between}={SS_{between}\over df_{between}}={128.944-29.833\over 2}=49.556"

"MS_{within}={SS_{within}\over df_{within}}={29.833\over 15}=1.989"

The F-statistic is computed as follows:


"F=\dfrac{MS_{between}}{MS_{within}}=\dfrac{49.556}{1.989}=24.916"

The following null and alternative hypotheses need to be tested:

"H_0:\mu_1=\mu_2=\mu_3"

"H_1:" Not all means are equal

The above hypotheses will be tested using an F-ratio for a One-Way ANOVA.

Based on the information provided, the significance level is "\alpha=0.01," and the degrees of freedom are "df_1=2, df_2=2," therefore, the rejection region for this F-test is "R=\{F: F>F_C=6.359\}"


Since it is observed that "F=24.916>6.359=F_C," it is then concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that not all 3 population means are equal, at the "\alpha=0.01"  significance level.


Using the P-value approach: The p-value is "p=0," and since "p=0<0.01," it is concluded that the null hypothesis is rejected. Therefore, there is enough evidence to claim that not all 3 population means are equal, at the "\alpha=0.01" significance level.